OpenStudy (kayc):

When f(x) is divided by ( x-3 ) and ( 2x+1 ), the remainders are 6 and -1 respectively. Find the remainder when f(x) is divided by ( 2x^2-5x-3 ).

OpenStudy (kayc):

Is the answer possibly 2x + 0 ...?

OpenStudy (raden):

first, 2x^2-5x-3 =(2x+1)(x-3), and suppose ax+b is remainders of f(x) when divided by 2x^2-5x-3. f(x) is divided by ( x-3 ) and its remainders is 6, it means f(3)=6 or 3a+b=3. ...(1) f(x) is divided by (2x+1 ) and its remainders is -1, it means f(-1/2)=-1 or -1/2*a+b=-1. ...(2) from (1) and (2), using elimination and subtitution method for solve a and b.

OpenStudy (raden):

yea, i got a=2, and b=0 so, its remainder is 2x+0=2x

OpenStudy (kayc):

Okay, thanks !

OpenStudy (raden):

welcome